[Brahms-l] Shift report 20050531 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Tue May 31 2005 - 08:00:58 EDT
Shifter leader was rk
Shifters were 


1. Very stable store all night, injected at about 22.30, is scheduled to be dumped at 09.00.
2. Runs: 15145 - 15156 (318k of tr.2 and 586k of tr.3)
3. Had problems with DAQ at the beginning of the shift - number of events didn't increment - found out it was problem with T2 VME crate - no problems since reseting.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue May 31 08:01:20 2005

This archive was generated by hypermail 2.1.8 : Tue May 31 2005 - 08:01:32 EDT